OpenStudy (anonymous):

While you are sitting on a park bench with your friend, you grab his hat and start running in a straight line away from him. Over the first 50m, you accelerate at 1.00m/s^2 up to your maximum running speed. You then continue at you maximum running speed for 5s more before your friend catches you. How far from the bench did you get before being caught?

OpenStudy (anonymous):

Strategy: 1) Find distance traveled when he accelerate. 2) Find distance traveled when he run with constant speed.

OpenStudy (anonymous):

1) x=v0t+1/2 * at^2 2) x=v*t

OpenStudy (anonymous):

ok , let me quickly try that out!

OpenStudy (anonymous):

the initial velocity is 0?

OpenStudy (anonymous):

Yep, he was sitting back then

OpenStudy (anonymous):

d= 0 + 1/2 (1m/s^2) t what do i do for the time? the question says 1m/s^2 to his max acce, but i dont know what is his maximum acceleration

OpenStudy (anonymous):

Heheh, actually, you can use v^2=v0^2+2ax to find v, (no need to use eq (1) then).

OpenStudy (anonymous):

then x= 50m?

OpenStudy (anonymous):

but arent u trying to find x? (How far from the bench did you get before being caught? )

OpenStudy (anonymous):

You already find x for first part (50 m). And we can use this x to vind his max velocity

OpenStudy (anonymous):

so he get caught at x=50..?

OpenStudy (anonymous):

No, he caught after running 5 sec with max velocity beyond that 50 m. So, find (2) first. He will be caught at distance 50 m + distance traveled at (2).

OpenStudy (anonymous):

but how can i do it if i dont know the maximum velocity

OpenStudy (anonymous):

x= v * t. i dont know both x and v

OpenStudy (anonymous):

Use this v^2=v0^2+2ax

OpenStudy (anonymous):

but i still dont know both x and v

OpenStudy (anonymous):

if i'm solving for (2) first, that means i 'm not using 50m

OpenStudy (anonymous):

x is 50 m here v0=0, and v is vmax you're looking for.

OpenStudy (anonymous):

and with the vmax, i sub it in x= v*t to find the distance after 50m?

OpenStudy (anonymous):

Yep, you're on it.

OpenStudy (anonymous):

this thing is so confusing..how can u guys even get this ==..omg..

OpenStudy (anonymous):

for the time, do i use d= v1 t + 1/2 a t ?

OpenStudy (anonymous):

Which time do you mean?

OpenStudy (anonymous):

How long did it take for your friend to catch you?

OpenStudy (anonymous):

You already know the time at second part (5 s). Now find for the first. Yes, you can use that for known a (1.00 m/s^2) and d (50 m). Alternatively v=v0+at (v and v0 also known)

OpenStudy (anonymous):

so if i find t using v=v0+at , do i need to add it to that 5s too?

OpenStudy (anonymous):

i got like over 100s, dont think that's right

OpenStudy (anonymous):

i got the right ans..dont know how and why i get it..but sure..thanks !

OpenStudy (anonymous):

Heheh, practice and you'll understand.